# A Multi-Perspective Analysis of Memorization in Large Language Models

### v3

@@ 4.4.1 Embedding Dynamics
In this section, we discuss the embedding of generated tokens for sentences memorized at different extents. We collect the hidden state at the last layer of each token. We compute the pair-wise Euclidean distance and Cosine similarity between sentences memorized at different extents across each generation step. Then, we perform dimension reduction on the embedding and visualize it in the 2D space using PCA.
Figure 5: Embedding Dynamics of Across Different Model Size. Memorized Token Num $x$ means $x$ tokens generated to align with the true continuation.
From Figure 5, we can obtain the following points [^2] Detailed numbers of Cosine Similarity and Euclidean distance between sentences with different memorization scores are at different decoding steps are at Appendix B.2:
@@
  • Additionally, we can see that for highly memorized sentences, even when they are not totally memorized. They are still close in the embedding space, indicating that the generated content may be both semantically and lexically similar to the memorized content.
  • Additionally, we can see that the large-size model has a higher Euclidean distance and lower Consine Similarity. The high Euclidean distance is caused by the expansion of hidden size (512 for 70m, 2048 for 1b, etc.). The decrease of the cosine similarity can also be attributed to the hidden size, as the increase in hidden size expands the embedding space. This gives more expressivity to the embedding space, as the Euclidean distance between sentence embedding also expands, which leads to the decrease of Cosine Similarity. This may also help to explain the performance gap between different model sizes. Large-size models have more expressivity in the embedding space; thus, different sentences with different meanings are more clearly distributed with fewer neighbors in the space. The smaller embedding space in a small model makes the mix of embedding easier, leading to ambiguity in the embedding space, thus degrading the performance.
@@ 4.4.2 Generation Dynamics and Entropy
In this section, we discuss how the model generates the memorized sentences with different memorization scores, e.g., the generation dynamics. As we have found the boundary effect in the input level, we also want to discuss the question of does the model exhibit certain behaviors when it generates memorized sentences or unmemorized sentences? To discuss this problem, we collect the generation probability at each token when the LLM generates memorized, unmemorized, and half-memorized sentences, calculate the entropy across 10,000 sentences, and plot the entropy at each token in Figure 6 and we can obtain the following from it:
  • First, we can see that the entropy is different depending on the extent of memorization. We can clearly see that an unmemorized sentence has a higher average entropy at each token than an unmemorized sentence. This shows that when the LLM generates unmemorized sentences, they are more unconfident about generating them.
  • Additionally, when the model starts to generate unmemorized tokens, we can see a significant increase in entropy, showing such boundary effect also exists in the decoding steps while behaving oppositely compared to the boundary effect in frequency analysis. We can also see a drop in entropy when decoding the memorized sequences. This corresponds to the fully memorized tokens and the first half of the half-memorized sequences. The entropy drops around the index of 32, which is the first generated token. However, we can see the drop of entropy for the memorized sentence is not samely significant as the half-memorized sentence, showing that even for the memorized content, the model is more confident when the sentence is fully memorized, leading to lower entropy.
  • We can also notice there is only one boundary at unmemorized or memorized sentences, further confirming that the model tends to memorize tokens continuously.
  • We can also see the entropy decrease with the increase in model size. This suggests that larger models are more confident about generating tokens than small ones. Additionally, we can see that the entropy is lower in the context of memorized tokens, showing that the entropy of context also relates to how a sentence is memorized.
@@
Table 2: Performance of Transformer Model on Prediction of Memorization. The context length is defaulted at 32. Length means the token length that is required to be predicted. Token and Full represent Token-Level Accuracy and Full Accuracy. The higher, the better. The best results across model sizes are bold, and the best results within a model size are underlined.
@@ 4.5.1 Results on Memorization of Prediction
  • Secondly, we showed that with the increase of the model size of LLM, the Token-level accuracy and Full-level accuracy increase rather than drop. This suggests that it is easier to predict the memorization behavior of larger-size models rather than small-size models as the embedding distance of large models is further, thus making the classification easier.
  • Thirdly, we found that with the increase in continuation size, the Token level accuracy gradually increases, possibly due to the increase in training data. For example, if the length to be predicted is $64$, the token used to train the model is four times higher than the prediction length of $16$. However, we see a drop in the Full accuracy as the increase in prediction length leads to more tokens being predicted. This makes reaching full accuracy harder as the tokens required to be predicted grow.
@@ 4.5.2 Analysis of Full Accuracy
In this experiment, we use the trained Transformer model to generate a case study for different memorized sentences. We presented the predicted ability for the corresponding token to be memorized. A higher probability means the token here is memorized, and a lower probability means the token here is unmemorized. The case study is presented in Figure 8. From the figure, we can tell:
  • For any model size, most sentences that are fully predicted share low memorization scores. The 00 score (unmemorized) sentence occupies the largest portion of fully predicted sentences. Meanwhile, we can see that sentences with high memorization scores are hard to predict fully and correctly. We can see that the $1$ score (memorized) sentence has the lowest share among fully predicted sentences, showing that memorized sentences are hard to predict.
  • Across model size, we can see that with the increase in model size, the proportion for unmemorized sentences increases. This suggests that the prediction model in the larger model is more accurate in predicting unmemorized sentences rather than memorized sentences.
  • The reason why a sentence with a low memorization score can be easily predicted can be due to the difference of significance at the boundary. As obtained from our previous experiments, the boundary effect of unmemorized sequences is more significant than memorized sequences in either token frequency or entropy of the generated token. Especially we can see that the boundary effect of memorized sequences decreases with model size while increasing for unmemorized sequences in previous results. This helps to explain why predicting memorization is easier in small-size models but harder in large-size models because the boundary effect becomes insignificant when we increase the model size. Similarly, the increase of boundary effect for unmemorized sentences with model size makes them easier to predict.
@@ 4.6 Conclusion
In this research, we comprehensively discussed the memorization behavior of LLMs across different perspectives. From a statistical level, we analyzed the memorization behavior of LLMs and expanded the analysis to sentences with lower memorization scores. Further transition analysis revealed the transition dynamics when trained in a large model, showing that most sentences stay in the original state while the memorized state is hard to transition to the unmemorized state. At the input level, through frequency analysis, we found the positive and negative boundary effects when generating memorized and unmemorized tokens. We showed the LLM tends to memorize longer sequences rather than short pieces. In the output decoding level, in the embedding level, we found clusters of sentences with different memorization scores in the embedding space, and the close distance of sentences with high memorization scores indicates the existence of paraphrase memorization. We analyzed the distribution level and calculated the entropy at the decoding dynamics. We found that when the LLM starts to generate an unmemorized sentence, the entropy boundary effect also exists and significantly changes with the model size and the memorization score of the corresponding sentence. Finally, we trained a Transformer model to predict the memorization. We found that prediction at the token level is easy, while prediction at the sentence level is challenging. Through analysis of correctly fully predicted samples, we found unmemorized tokens are easy to predict while memorized tokens are hard to predict, and such a phenomenon can be explained by the significance of the boundary effect.
@@ B.2 Supplementary Data for Embedding Dynamics
Similarly to Figure 5, we also provide detailed cosine similarity and distance results from Figure 9 to Figure 17. From those figures, we can see that the cosine similarity fluctuates but remains relatively stable for sentences with different memorization scores. Additionally, sentences that are not exactly the same but close regarding the memorization score are very close in the embedding space. For example, the fully memorized sentence is also close to sentences with high memorization scores with embedding similarity over 0.9. This shows the possibility that the model is generating paraphrased memorized sequences. However, with the increase in the model size, the cosine similarity decreases. For the Euclidean distance, we can see that the embedding distances have a decreasing trend with the increasing decoding steps, while the mutual Euclidean increases with the model size.

### v4

@@ 4.4.1 Embedding Dynamics
In this section, we discuss the embedding of generated tokens for sentences with different memorization scores. We collect the hidden state at the last layer of each generated token for sentences with different memorization scores and compute the pair-wise Euclidean distance and cosine similarity between them. Then, we perform PCA F.R.S. (1901) on those embeddings and visualize as shown in Figure 5, and we can obtain:
Figure 5: Embedding Dynamics of Across Different Model Size. Memorized Token Num $x$ means $x$ generated tokens are the same with the true continuation.
 [^3] Detailed numbers of Cosine Similarity and Euclidean distance between sentences with different memorization scores are at different decoding steps are at Appendix B.2
  1. (I)
@@
  2. (II)
    For highly memorized sentences, even when not fully memorized, they cluster closely in the embedding space. This suggests the generated content is both semantically and lexically similar to the memorized content, indicating the existence of paraphrased memorization.
  3. (III)
    Larger models exhibit higher Euclidean distances and lower cosine similarities. The higher Euclidean distance is due to the expansion of hidden sizes (e.g., 512 for 70 million, 2048 for 1 billion parameters). The decrease in cosine similarity is also attributed to the expanded hidden sizes, which increase the expressivity of the embedding space. This expansion enlarges the sentence embeddings, reducing cosine similarity but enhancing the differentiation between sentences. This expanded expressivity in embedding helps to explain the performance gap between different model sizes: larger models distribute different sentences more distinctly with fewer embedding overlaps, while smaller models mix embeddings more, leading to ambiguity and degraded performance.
@@ 4.4.2 Generation Dynamics and Entropy
This section discusses how the model generates the memorized sentences with different memorization scores, e.g., the generation dynamics. Given the observed boundary effect at the input level, we aim to address whether the model exhibits specific behaviors when generating memorized versus unmemorized sentences. To investigate this, we collected the generation probability for each token during the generation of memorized, unmemorized, and half-memorized sentences. We then calculated the entropy across 10,000 sentences and plotted the entropy for each token in Figure 6. From this analysis, we derive the following observations:
  1. (I)
    First, we can see that the entropy differs depending on the extent of memorization and model size. An unmemorized sentence has a higher average entropy at each token than an unmemorized sentence. This shows LLM is more confident when generating memorized content.
  2. (II)
    Additionally, when the model starts to generate unmemorized tokens, we can see a significant increase in entropy, showing such boundary effect also exists in the decoding steps while behaving oppositely compared to the boundary effect in frequency analysis. The entropy drops around the index of 32 for the memorized and half-memorized content. This also shows the existence of the boundary effect for memorized content. However, we can see the drop of entropy for the memorized sentence is not samely significant as the half-memorized sentence, showing that even for the memorized content, the model is more confident when the sentence is fully memorized, leading to lower entropy.
  3. (III)
    We can also see the entropy decrease with the increase in model size. This suggests that larger models are more confident about generating tokens than small ones. Additionally, we can see that the entropy is lower in the context of memorized tokens, showing that the entropy of context also relates to how a sentence is memorized. Further, the significance of the boundary effect in entropy decreases with the model size for memorized content while remaining unchanged for unmemorized content.
@@
Table 2: Performance of Transformer Model on Prediction of Memorization. The context length is defaulted at 32. Length means the token length that is required to be predicted. Token and Full represent Token-Level Accuracy and Full Accuracy. The best results across model sizes are bold, and the best results within a model size are underlined.
@@ 4.5.1 Results on Memorization of Prediction
  2. (II)
    As the model size of LLMs increases, both token-level and full-level accuracy improve. This suggests that predicting memorization behavior is easier for larger models because the greater embedding distances make classification easier.
  3. (III)
    Token-level accuracy increases with the continuation size, likely due to increased training data. For instance, a prediction length of 64 involves four times more training data than a prediction length of 16. However, full-sentence accuracy decreases as the prediction length increases, making it harder to achieve full accuracy because more tokens need to be predicted correctly.
@@ 4.5.2 Analysis of Full Accuracy
In this experiment, we analyze how fully correct predictions distribute across memorization scores to discuss whether the difficulty in predicting memorization changes with the memorization score. From the results shown in Figure 7, we can tell:
  1. (I)
    For any model size, they are more good at predicting sentences with low memorization scores. In contrast, the low portion for sentences with high memorization scores indicate they are harder to predict accurately.
  2. (II)
    As model size increases, the proportion of low memorization scores raise and decrease for high memorization score sentences which even reaches zero for 6.9b model. This suggests larger models are more accurate in predicting unmemorized sentences compared to memorized ones.
  3. (III)
    The ease of predicting sentences with low memorization scores can be attributed to the significance of the boundary effect. Previous experiments show that the boundary effect of unmemorized sequences is more obvious than that of memorized sequences in both token frequency and entropy. Particularly, the boundary effect decreases for memorized sequences while increases for unmemorized sequences with increased model size. Such change makes the boundary effect less significant for memorized sentences, making them harder to predict, while more significant for unmemorized sentences, making them easier to predict.
@@ 4.6 Conclusion
In this study, we comprehensively discussed LLM memorization from various perspectives. From a statistical level, our analysis extended the scope of previous research to sentences with lower memorization scores. This is followed by an experiment showing the memorization transition across model sizes. At the input dynamic level, through frequency analysis, we found the positive and negative boundary effects when generating memorized and unmemorized tokens and their relation to how easily a sentence can be memorized/unmemorized. In the output dynamics, at the embedding level, we found clusters of sentences with different memorization scores in the embedding space, and the close distance of sentences with high memorization scores indicates the existence of paraphrase memorization. At the entropy level, we observed an opposite boundary effect and analyzed its change with model size. Finally, we trained a Transformer model to predict the memorization, showing that token-level prediction is easy while sentence-level is challenging. Through analysis of fully correct predicted samples, we found unmemorized tokens are easier to predict than memorized tokens, which can be explained by the significance of the boundary effect.
@@ B.2 Supplementary Data for Embedding Dynamics
Similarly to Figure 5, we also provide detailed cosine similarity and distance results from Figure 9 to Figure 17. From those figures, we can see that the cosine similarity fluctuates but remains relatively stable for sentences with different memorization scores. Additionally, sentences that are not exactly the same but close regarding the memorization score are very close in the embedding space. For example, fully memorized sentences are also close to sentences with high memorization scores and embedding similarity of over 0.9. This shows the possibility that the model is generating paraphrased memorized sequences. However, with the increase in the model size, the cosine similarity decreases. For the Euclidean distance, we can see that the embedding distances have a decreasing trend with the increasing decoding steps, while the mutual Euclidean increases with the model size.